OpenStudy (anonymous):

Find a simple transformation of g(x) that is identical to f(x) for all points except those indicated?? f(x)= 2x^2+5x+2 / x^2+2x; g(x)=1/x; x=-2

OpenStudy (anonymous):

So you mean identical to f(x) except for x=-2?

OpenStudy (anonymous):

the problem is that your function is not defined at x = -2

OpenStudy (anonymous):

yea?

OpenStudy (anonymous):

so factor the numerator and denominator, cancel, and you will get your new function

OpenStudy (anonymous):

\[f(x)=\frac{2x^2+5x+2}{x^2+2x}=\frac{(2x+1)(x+2)}{x(x+2)}=\frac{2x+1}{x}\]

OpenStudy (anonymous):

is that the same as 2+1/x? cause thats the answer?

OpenStudy (anonymous):

wow site is having a fit. can you read what i wrote?

OpenStudy (anonymous):

yea i can

OpenStudy (anonymous):

yes if you divide both terms by x you get \[2+\frac{1}{x}\]

OpenStudy (anonymous):

idea is clear right? your function is not defined at -2 so you factor and cancel and get another function that is the same, but is defined at -2

OpenStudy (anonymous):

ohkay gotchya yea it is clear! could you help me with just one more? Im just not sure how to simplify this one?

OpenStudy (anonymous):

sure post and i will be right there